1000-kg sports car accelerates from rest to 95 km/
Answer #1

Given,

Mass, m= 1000 kg

Velocity, v= 95 km/h = 26.4 m/s

(1 km/h = (5 / 18) m/s)

\therefore Energy, E = (1 / 2) mv2

E = (1 / 2) (1000) (26.4)2 J

E = 348480 J

Now, Power, P = (Energy / Time) = (348480 / 6) J/s

P = 58080 W